### v1

# The Generation Phases of Flow Matching: a Denoising Perspective

Abstract
Flow matching has achieved remarkable success, yet the factors influencing the quality of its generation process remain poorly understood. In this work, we adopt a denoising perspective and design a framework to empirically probe the generation process. Laying down the formal connections between flow matching models and denoisers, we provide a common ground to compare their performances on generation and denoising. This enables the design of principled and controlled perturbations to influence sample generation: noise and drift. This leads to new insights on the distinct dynamical phases of the generative process, enabling us to precisely characterize at which stage of the generative process denoisers succeed or fail and why this matters.
1 Introduction
Flow matching (FM, Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023) and diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) have achieved state-of-the-art results in generating images, videos, audio, and even text, where they are able to produce content that is virtually indistinguishable from human-produced ones. Research in the field remains extremely active, with many important questions still open: improving sample quality, making training and inference more efficient (Karras et al., 2022; Rombach et al., 2022) and, perhaps most importantly, understanding why current generative models perform so well.
Despite their striking successes, the precise mechanisms that make current generative models so effective still remain elusive. Identifying those is critical to improve them, and several leads have been proposed: analyzing the behaviour of the generative process across time (Biroli et al., 2024), exploiting connections with the exact minimizer of the training loss (Kamb and Ganguli, 2025; Niedoba et al., 2025), explaining memorisation and generalisation (Kadkhodaie et al., 2024; Sclocchi et al., 2025), or the impact of optimization procedures (Wu et al., 2025). Some of these studies, of theoretical nature, have provided elegant interpretations (e.g. the “target stochasticity” of Vastola, 2025, to explain generalisation), but were later questioned by empirical findings (Bertrand et al., 2025). This highlights the need for carefully designed empirical frameworks that can probe such theories and, in doing so, guide the development of better methods through deeper theoretical understanding.
In this work, we aim to bring new elements of understanding to the behaviour of flow matching by exploiting a denoising perspective. To this end, we construct a toolkit of denoisers, which differ in their parametrizations and the weighting schemes applied to their training losses. This perspective allows us to directly relate denoising performance to generative performance. By leveraging the equivalence between learning the ideal velocity field in flow matching and learning an ideal denoiser at each time step, we use this toolkit to address the following questions:
Is a good generative model essentially nothing more than a good denoiser at every noise level?
Are there specific times during the generative process where accuracy matters most?
How do early versus late phases of generation contribute to generalisation and sample quality?
To answer these questions, our contributions are:
  1. 1.
    We design a denoising toolkit, namely, controlled procedures to test the impact of several factors on the performance of flow matching models. We show that different denoising losses and parametrizations, though theoretically equivalent if perfectly trained, lead to very different empirical performance, where denoising and generation quality are strongly related.
  2. 2.
    We engineer two types of controlled perturbations applied on the generation process: drift- and noise-type perturbations. We show that they impact distinct temporal phases of generation. We further analyze these generation stages by exhibiting a discrepancy between the spatial regularities of target and learned velocity fields, at early times.
  3. 3.
    We exploit the generality of the denoising framework, that allows building new models in a principled manner, with similar generation quality (FID-wise) but different generation behaviours. We highlight the importance of the intermediate stage for generation by learned models, a stage that is not revealed by the closed-form optimal velocity.
The structure of the paper is the following: Section 2 provides an introduction to flow matching; Section 3 lays down the theoretical equivalence between flow matching and denoising; Section 4 is dedicated to related works. Section 5 provides an empirical probe of the theoretical equivalence between denoising and generation. Section 6.1 leverages our framework to design relevant modifications of the generative process. Finally, Section 6 provides new insights on the phases of generation and their nature.
2 Background on flow matching
The generative process is defined over time $t\in[0,1]$, with an initial sample $x_{0}\sim p_{0}$ and a target sample $x_{1}\sim p_{1}$. To connect with the concept of denoisers in the sequel, we further assume that the latent distribution is standard Gaussian: $p_{0}=\mathcal{N}(0,\mathrm{I}_{d})$, and we work in the setting where the coupling $p_{(x_{0},x_{1})}$ is the product coupling $p_{0}\otimes p_{1}$. In flow matching, generation of new samples is performed via the numerical resolution of an ordinary differential equation (ODE) on $[0,1]$:
[…]
the function $v:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ being called the velocity. The generated sample is simply the ODE solution at time $t=1$, namely $x(1)$; for an appropriate velocity, it should behave like a sample from $p_{1}$. In practice, the velocity is parametrized by a neural network $v_{\theta}$ and learned by solving:
$\min_{\theta}\,\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ x_{0}\sim p_{0},\\ x_{1}\sim p_{1}\end{subarray}}\left[\|v_{\theta}(x_{t},t)-(x_{1}-x_{0})\|^{2}\right],$ (2)
where $x_{t}:=(1-t)x_{0}+tx_{1}$ is the linear interpolation between $x_{0}$ and $x_{1}$. It is well-known that the solution $v^{\star}$ to this minimization problem (over all measurable functions) is given by a conditional expectation:
$v^{\star}(x_{t},t)=\mathbb{E}[x_{1}-x_{0}\mid x_{t},t].$ (3)
In practice, sampling from $p_{1}$ in (2) is impossible, and points $x_{1}$ are instead drawn from a dataset $x^{(1)},\ldots,x^{(n)}$ of samples from $p_{1}$. Effectively, $p_{1}$ is replaced in (2) by the empirical measure $\hat{p}_{1}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x^{(i)}}$. This has an important consequence: the minimizer of (2), when $p_{1}$ is replaced by $\hat{p}_{1}$, admits a closed-form $\hat{v}^{\star}$ (Gu et al., 2025; Bertrand et al., 2025):
$\hat{v}^{\star}(x,t)=\sum_{i=1}^{n}\lambda_{i}(x,t)\frac{x^{(i)}-x}{1-t},\quad\text{where}\quad\lambda_{i}(x,t)=\operatorname{softmax}\left(\left(-\tfrac{\|x-tx^{(j)}\|^{2}}{2(1-t)^{2}}\right)_{j}\right)_{i}.$ (4)
Paradoxically, using this closed-form velocity for generation can only reproduce samples from the training set. At the same time, this target velocity exhibits different behaviours across time: at $t=0$, it points towards the dataset mean, while as $t\to 1$ it points towards a single training point. Thus, understanding why trained FM models work comes down to determining, qualitatively and quantitatively, at which times $t$ and points $x$ the models must deviate from the closed-form in order to generate new samples, still consistent with the data distribution.
3 Equivalence between flow matching and denoising
3.1 From flow matching to denoising
We now lay down the procedure to construct a denoiser from a flow matching model. Using the identity $x_{t}=(1-t)x_{0}+tx_{1}$ and the expression of the optimal velocity field (3), ones obtains the denoising identity111the denoiser is interchangeably written $D(x,t)$ or $D_{t}(x)$ for concision
$D_{t}^{\star}(x_{t}):=\mathbb{E}[x_{1}|x_{t},t]=x_{t}+(1-t)v^{\star}(x_{t},t),$ (5)
namely the Minimum Mean Square Error estimator of the clean image $x_{1}$ given the noisy observation $x_{t}$ at time $t$. Thus, any optimally trained FM model naturally yields an optimal denoiser. Building on (5), several works on flow-matching for image restoration define a denoiser at time $t$ as $D_{t}:x_{t}\mapsto x_{t}+(1-t)\,v(x_{t},t)$, see e.g. Zhang et al. (2024); Pokle et al. (2024); Martin et al. (2025).
Figure 1: Equivalence between velocity $v_{t}$ and denoiser $D_{t}$. Learning the optimal velocity amounts to learning an optimal denoiser at every time $t$.
3.2 From denoisers to velocities: the denoising toolkit
The same way a velocity field can be mapped to its associated denoiser, any denoiser $D$ induces a velocity field $v(x,t)=\frac{D_{t}(x)-{\rm Id}}{1-t}$. This duality yields the question at the heart of our study: is flow matching nothing more than learning a denoiser at all possible noise levels, and then sampling by following the velocity derived from it?
To investigate this, we propose a systematic way of constructing denoisers. Throughout the paper, we use the term denoiser in a broad sense: a function that maps a noisy input, obtained by corrupting $x\sim p_{1}$, to a clean estimate of $x$. We will consider generative denoisers $D_{t}$, defined for each $t\in[0,1]$ and taking as input images of the form $x_{t}=tx+(1-t)\varepsilon$, $\varepsilon\sim\mathcal{N}(0,I_{d})$.
Remark 1 (Equivalence of generative and classical denoisers).
Generative denoisers differ from classical denoisers $D_{\sigma}$, parameterized by a noise level $\sigma$ and taking inputs of the form $x_{\sigma}=x+\sigma\varepsilon$. These two forms are equivalent up to rescaling. If the denoiser is parameterized by a noise level but the input is given as $x_{t}=tx+(1-t)\varepsilon$, one can define $x_{\sigma}=\frac{x_{t}}{t}$ and apply the classical denoiser $D_{\sigma}$ with $\sigma=\frac{1-t}{t}$ to approximately recover $x$. Conversely, if the input is of the form $x_{\sigma}=x+\sigma\varepsilon$, setting $x_{t}=\frac{x_{\sigma}}{1+\sigma}$ and $t=\frac{1}{1+\sigma}$, a generative denoiser $D_{t}$ can be applied to $x_{t}$ to recover $x$. The mapping from $\sigma$ to $t$ is a bijection between $0,\infty)$ and $(0,1]$.
Equipped with this definition, we introduce a denoising toolkit, a family of neural denoisers obtained by numerically solving optimization problems of the form
$\boxed{\underset{D\in\mathcal{C}}{\mathrm{minimize}}~\mathcal{L}(D),}$ (6)
where $\mathcal{C}$ is the class of functions parameterized by a neural network, and $\mathcal{L}(D)$ is the training loss. For specific $\mathcal{C}$ and $\mathcal{L}$, we recover the standard flow matching model, but making these design choices explicit and decoupling them, this abstraction opens the way to a systematic study of their impact.
3.3 Denoising losses $\mathcal{L}$
We present three types of denoising losses, all expressed for a generative denoiser $D:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$, taking as input $x_{t}=(1-t)x_{0}+tx_{1}$, with clean image $x_{1}\sim p_{1}$ and noise $x_{0}\sim\mathcal{N}(0,I_{d})$. Detailed derivations are provided in Appendix A.
Flow matching denoising loss. The first loss arises directly from the flow matching objective in Eq. (2). Substituting the velocity $v(x,t)=\tfrac{D(x,t)-x}{1-t}$ into Eq. (2) together with $x_{1}-x_{0}=\frac{x_{1}-x_{t}}{1-t}$, one can write the standard FM objective as a function of the denoiser,
$\mathcal{L}_{\text{FM}}(D):=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ x_{0}\sim\mathcal{N}(0,I_{d})\\ x_{1}\sim p_{1}\end{subarray}}\left[\frac{1}{(1-t)^{2}}\|D(x_{t},t)-x_{1}\|^{2}\right].$ (7)
Thus, training a denoiser under FM amounts to minimizing a weighted Mean Squared Error (MSE), where the error at time $t$ is weighted by $w_{t}^{\text{FM}}:=(1-t)^{-2}$.
Classical denoising loss. As we have seen, classical denoisers are parameterized by a noise level $\sigma$ and trained on inputs $x_{\sigma}=x_{1}+\sigma x_{0}$. Such a denoiser $\tilde{D}$ is usually trained on noise levels ranging from 0 to $\sigma_{\max}$, by minimizing
$\mathcal{L}(\tilde{D})=\mathbb{E}_{\begin{subarray}{c}\sigma\sim\mathcal{U}([0,\sigma_{\max}])\\ x_{0}\sim\mathcal{N}(0,I_{d})\\ x_{1}\sim p_{1}\end{subarray}}\left[\|\tilde{D}(x_{\sigma},\sigma)-x_{1}\|^{2}\right].$ (8)
Using the equivalence between the $\sigma$- and $t$-parameterizations (Remark 1) and the change of variables $\sigma=\tfrac{1-t}{t}$, one can rewrite this loss as
$\mathcal{L}_{\text{classic}}(D):=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([1/(1+\sigma_{\max}),1])\\ x_{0}\sim\mathcal{N}(0,I_{d})\\ x_{1}\sim p_{1}\end{subarray}}\left[\frac{1}{t^{2}}\|D(x_{t},t)-x_{1}\|^{2}\right].$ (9)
Compared to the FM loss $\mathcal{L}_{\text{FM}}$, classical denoising therefore differs in two important ways: (i) the loss includes a weight $w_{t}^{\text{classic}}:=\mathbf{1}_{[(1+\sigma_{\max})^{-1},1]}(t)\cdot t^{-2}$; (ii) the range of $t$ is truncated to $1/(1+\sigma_{\max}),1]$, which covers the full interval $[0,1]$ only if $\sigma_{\max}=\infty$. In other words, classical denoisers cannot handle small times in FM except if trained with unbounded noise levels. In practice, we set $\sigma_{\max}=19$ so that [^2] In traditional denoising, models are usually trained with noise level at most $\sigma=100/255\simeq 0.4$. $t_{\text{min}}=1/(1+\sigma_{\max})=0.05$.
Unweighted denoising loss. A natural comparison baseline is to use “no weights” and train with the plain mean squared error (i.e., weighting $w_{t}^{\text{den}}:=1$):
$\mathcal{L}_{\text{den}}(D):=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ x_{0}\sim\mathcal{N}(0,I_{d})\\ x_{1}\sim p_{1}\end{subarray}}\left[\|D(x_{t},t)-x_{1}\|^{2}\right].$ (10)
More general weightings in denoising losses. The three losses considered above emphasize opposite time intervals: $\mathcal{L}_{\mathrm{FM}}$ stresses large $t$ (lightly corrupted inputs), while $\mathcal{L}_{\text{classic}}$ stresses small $t$ (highly corrupted inputs). This is expressed by the time-dependent weightings $w_{t}$ in the loss $\mathbb{E}_{t,x_{0},x_{1}}\left[w_{t}\|D(x_{t},t)-x_{1}\|^{2}\right].$ As part of our methodology, in Section 6 we will also explore handcrafted weightings, allowing us to probe the generation process.
3.4 Equivalence in the optimal setting
A critical point is that all the above losses share the same minimizer in $L^{2}(\mathbb{R}^{d}\times[0,1])$, namely the MMSE denoiser $D^{\star}(x_{t},t)=\mathbb{E}[x_{1}|x_{t},t]$. However, in practice, minimization is restricted to a parametric function class $\mathcal{C}$ parametrized by a given neural architecture, optimization algorithms may be imperfect, and thus different loss weightings $w_{t}$ can lead to different numerical solutions.
3.5 Parametrizations of the denoisers and velocities
Table 1: Summary of the denoising losses (left) and parametrization classes (right).
Losses $\mathcal{L}$ Parametrization classes $\mathcal{C}$
$\mathcal{L}_{\text{FM}}:\,w_{t}^{\text{FM}}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{NN}}$: $D(x,t)=N^{\theta}(x,t)$
$\mathcal{L}_{\text{classic}}:\,w_{t}^{\text{classic}}=\mathbf{1}_{[(1+\sigma_{\max})^{-1},1]}(t)\cdot t^{-2}$ $\mathcal{C}_{\mathrm{I+NN}}$: $D(x,t)=x+(1-t)N^{\theta}(x,t)$
$\mathcal{L}_{\text{den}}:\,w_{t}^{\text{den}}=1$
We now describe two parametrization classes $\mathcal{C}$ for the denoisers, which are used in the minimization of the losses $\mathcal{L}$ defined in Section 3.3. In all cases, $N^{\theta}$ denotes a neural network, with parameters $\theta$ belonging to some set $\Theta$.
Class $\mathcal{C}_{\mathrm{NN}}$: Standard neural network parametrization. A straightforward approach is to directly parametrize $D$ by a neural net $N^{\theta}$ taking as input the noisy image and the time:
$\mathcal{C}_{\mathrm{NN}}=\left\{D:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}\ \middle|\ D:x,t\mapsto N^{\theta}(x,t),\,\theta\in\Theta\right\}.$ (11)
Class $\mathcal{C}_{\mathrm{I+NN}}$: Residual denoiser form $D=\mathrm{Id}+(1-t)N^{\theta}$. Following the relationship between the optimal FM denoiser $D^{\star}$ and the optimal velocity field $v^{\star}$, namely $D^{\star}={\rm Id}+(1-t)v^{\star}$, one can also parametrize the denoiser in residual form, where the network acts as a correction to identity:
$\mathcal{C}_{\mathrm{I+NN}}=\left\{D:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}\ \middle|\ D:x,t\mapsto x+(1-t)N^{\theta}(x,t),\,\theta\in\Theta\right\}.$ (12)
A key feature of this parametrization is that it enforces at $t=1$, $D_{t}={\rm Id}$, meaning the denoiser leaves its input unchanged which is the expected behavior (no noise in the input). It also matches the parametrization of standard flow matching models, where the velocity field is directly parametrized by a neural network. The different losses and parametrizations considered are summarized in Table 1.
4 Related works
Contrary to most previous works, we do not seek to further refine the loss function or propose new parameterizations. Instead, we take the opposite approach: rather than hypothesizing which timesteps are most important and designing new losses accordingly, we fix several existing weighting schemes and systematically analyze their impact on generation behavior. Our goal is twofold: (i) to determine whether the generation process can be interpreted as a form of classical denoising operating across a broader range of noise levels, and (ii) to dissect the distinct temporal phases that occur during generation.
Weighting strategies in denoising score matching.
We now discuss related works within the diffusion framework, which is known to be equivalent to flow matching when the source distribution $p_{0}$ is Gaussian (Gao et al., 2025). For simplicity, we consider the variance-preserving diffusion process where $x_{t}$ is obtained from a clean sample $x_{0}$ and a Gaussian noise $\varepsilon\sim\mathcal{N}(0,I_{d})$ as $x_{t}:=\alpha_{t}x_{0}+\sigma_{t}\varepsilon$ with $\alpha_{t}^{2}+\sigma_{t}^{2}=1$. Note that we use the standard diffusion notations in this section only, meaning that the sampling process evolves from $t=T$ (noise) to $t=0$ (clean image), as opposed to the flow matching convention where $t$ evolves from 0 to $1$. Regarding parametrization, most diffusion implementations train a network to predict the noise $\varepsilon$. Some papers (Salimans and Ho, 2022; Hang et al., 2023) also study $v$-prediction (i.e. $\mathcal{C}_{\mathrm{I+NN}}$) or $x$-prediction (i.e. $\mathcal{C}_{\mathrm{NN}}$).
Regarding weighting, most diffusion papers follow the formulation from Ho et al. (2020) which uses an unweighted loss in the $\varepsilon$-prediction, i.e. $\mathbb{E}_{t,x_{0},\varepsilon}[\|\varepsilon-\varepsilon^{\theta}(x_{t},t)\|^{2}]$, corresponding in the denoising framework to the weighting ${\alpha_{t}^{2}}/{\sigma_{t}^{2}}$, equal to the signal-to-noise ratio. A line of work is devoted to crafting loss weightings. Salimans and Ho (2022) propose the weighting $\max(\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}},1)$ while Yu et al. (2024) use weighting $\frac{\alpha_{t}}{\sigma_{t}}$. Both approaches assign greater importance to large noise levels, arguing that these correspond to more difficult denoising tasks and play a crucial role in error propagation during sampling. Interpreting diffusion training as multitask optimization, Hang et al. (2023) observe conflict between different timesteps objectives and suggest weighting $\min(\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}},\gamma)$ to avoid putting too much weight on the low noise levels, identified as an easy denoising task. In contrast, Choi et al. (2022) introduce the P2 weighting that puts more weight on the intermediate times, hypothesizing that perceptual features emerge during this content phase, as opposed to the early coarse phase or the late cleanup phase, where little noise remains in the image.
Rather than hypothesizing which timesteps are important or directly relating the weighting strategy to FID performance, we introduce an intermediate test metric–the PSNR evaluated at each timestep–which enables a more fine-grained analysis of how different timesteps contribute to generation quality. Similar to Choi et al. (2022), we find that effective denoising at intermediate times is crucial for high-quality generation.
Unifying perspectives on loss weightings.
Kingma and Gao (2023) provide a unifying view on a wide range of loss weightings (including FM and the previous mentioned ones) by rewriting them as differently weighted ELBO formulations. Kumar et al. (2025) also lay down the different losses (noise prediction, score prediction, etc) and the loss weighting they induce.
Nevertheless, they works do not analyze why different weighting choices can affect FID performance. We address this question through our denoising toolkit.
Linking diffusion and denoising.
Beyond standard diffusion, Delbracio and Milanfar (2023) study a broader “degradation-to-clean” setting (which is not limited to Gaussian denoising) and pick a standard denoising loss (i.e. weighting $w_{t}^{\mathrm{den}}=1$, parametrization $\mathcal{C}_{\mathrm{NN}}$). For diffusion, they report good results on CelebA ($64\times 64$), yet below state-of-the-art levels, which is consistent with our observations. Leclaire et al. (2025) also provide a synthetic overview of the connections between diffusion models and classical additive Gaussian denoising, interpreting diffusion as an iterative “Noising-Relaxed Denoising” process to better understand the role of noise schedules.
Analysis of the generation phases.
Several works study the generation process to shed light on generalisation, i.e. why they are able to generate samples that do not belong to the training dataset. This question is closely linked to how trained models approximate the exact minimizer of the training loss, which admits a closed-form expression and which can only reproduce training samples.
One line of research focuses directly on the closed-form, either theoretically (Biroli et al., 2024) or empirically (Bertrand et al., 2025). Both works highlight a critical time, at early timesteps, beyond which the closed-form points towards a single training example, while the trained models begin to deviate from it.
Another group of works seeks to understand the reasons for generalisation by characterizing the velocities or scores that are actually learned. Kadkhodaie et al. (2024); Niedoba et al. (2025); Kamb and Ganguli (2025) analyze the effective receptive field of trained models and show that it evolves from global (at high noise levels) to local (at low noise levels). Niedoba et al. (2025); Kamb and Ganguli (2025) both argue that what is actually learned by the models is a patch-wise version of the closed-form, with patch size determined by the model’s receptive field: they demonstrate that this new formulation can predict in certain cases, without training, the samples learned by the model.
Beyond works focusing on the closed-form and its approximation, Sclocchi et al. (2025) connect the diffusion generative process to the learning of high- and low-level features: they specifically show that there exists a critical time at which image class is determined, while low-level features evolve smoothly throughout the generation process.
Finally, some works explain generalisation by imperfect or early-stopped optimization (Wu et al., 2025; Bonnaire et al., 2025; Favero et al., 2025).
In contrast to previous analyses that study the generation phases by measuring the deviation from the closed-form solution (based on the finite training data), we adopt a test-time denoising perspective. Rather than relying on a purely training-oriented metric, we evaluate the PSNR on test data to indirectly measure the deviation from the ideal MMSE denoiser $\mathbb{E}_{x_{1}\sim p_{1}}[x_{1}\mid x_{t},t]$. This approach provides a fine-grained, time-specific analysis of how well the model performs denoising at each timestep, and how this relates to overall generative quality (e.g., FID).
5 Generation: denoising at every time level?
(a) Difference in PSNR (lower is better) between standard FM ($\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$) and various models, computed on 1000 test images. Positive values indicate worse denoising performance compared to standard FM.
(b) FID on 10k tests images (lower is better).
Figure 2: PSNR and FID for the different losses and parametrizations, CIFAR-10. Models that reach the highest PSNR (low difference in PSNR compared to standard FM) also reach the lowest FID.
5.1 Denoising and generative metrics
As a first investigation, we evaluate models trained using the different couples of denoising losses/parametrizations ($\mathcal{L}$, $\mathcal{C}$) on CIFAR-10 ($\times32$, Krizhevsky and Hinton, 2009) and CelebA-64 ($\times64$, Yang et al., 2015). All models share the same architecture and training hyperparameters, borrowed from the standard FM training (full details in Appendix B). We also train 10 independent denoisers with $\mathcal{L}_{\mathrm{den}}$, each trained only on the time interval $[i/10,(i+1)/10]$ for $i\in\llbracket 0,9\rrbracket$, yielding a velocity field defined in a piecewise manner over time; the resulting model is denoted “10-denoisers”.
Figure 2 displays the performance of our trained models both in denoising and in generation. Denoising at noise level/time $t$ is evaluated using the Peak Signal-to-Noise Ratio (PSNR) between denoiser output $D_{t}(x_{t})$ and clean sample $x_{1}\sim p_{1}$, averaged on $1000$ test images. Generation is evaluated using the Fréchet Inception Distance (FID, Heusel et al., 2017) between 10k test samples and 10k generated samples. First, Figure 2 confirms that, although all denoising losses $\mathcal{L}$ are equivalent in terms of the optimal denoiser they promote, both the choices of $\mathcal{L}$ and $\mathcal{C}$ impact performance in practice [^3] We provide in Appendix C a table with additional tested weightings, showing the same trends.. For every loss except the $\mathcal{L}_{\text{classic}}$ (for which the generated images are of such poor quality that the corresponding FID values are not meaningful), the residual parametrization $\mathcal{C}_{\mathrm{I+NN}}$ consistently outperforms the plain parametrization $\mathcal{C}_{\text{NN}}$, supporting the hypothesis that explicitly enforcing $D_{t}={\rm Id}$ introduces a beneficial implicit bias. Second, the models with lower FID also have better PSNRs at all noise levels, except for 10-denoisers which has worse PSNR than standard flow matching for $t<0.9$. The best performance is obtained with the FM loss under the parametrization $\mathcal{C}_{\mathrm{I+NN}}$ (i.e. the standard flow matching approach). On the contrary, the models trained with the classical loss ($w_{t}^{\text{classic}}=t^{-2}$) and the unweighted denoising loss ($w_{t}^{\text{den}}=1$), both motivated only by denoising considerations, obtain not only poorer generation performance, but also poorer denoising performance. Combined with the good performance of the 10-denoisers, this suggests that training a single network (taking $t$ as a parameter) to denoise at every noise level, if not counterbalanced with an appropriate use of weights, is detrimental to performance. Perhaps surprisingly, the FM weights $w_{t}^{\text{FM}}=1/(1-t)^{2}$ turn out to be the most efficient for denoising, although they put more emphasis on accurate denoising at low noise level ($t$ close to $1$), which one may think of as an easy task.
On top of these general trend, the behaviour of 10-denoisers is more complex: although it yields worse PSNRs than the FM denoiser (except at $t\geq 0.9$), it still manage to reach a comparable FID.
5.2 Inpainting
We complement our experiments with a third metric, at the crossroads of denoising and generation, to confirm the correlation previously observed. To this end, we turn to an intermediate task: image inpainting. This task is both an inverse problem, a field in which denoisers are regularly used in a plug-and-play fashion (Venkatakrishnan et al., 2013; Meinhardt et al., 2017), and a generative challenge, as the model must synthesize large missing parts of the image. We evaluate our denoisers in this setting using the PnP-Flow algorithm (Martin et al., 2025), which incorporates FM models into plug-and-play frameworks. This method builds on the equivalence between denoisers and FM velocities, since in PnP-Flow the time-dependent denoiser is directly induced by the learned velocity field. Experimental details and visual examples are provided in Appendix E, Figure 9. As shown in Figure 3, the ranking of models on the inpainting task coincides with their ranking in terms of FID and PSNR. Now that it is clear that the considered models are uniformly good or bad across all three metrics, we investigate in more depth the factors that may cause this performance gap.
Figure 3: Inpainting results in terms of PSNR (higher is better) as a function of the time in PnP-Flow, CelebA-64. Results are averaged over 100 images. Mask of size $17\times 17$. The horizontal black line represents as a reference the PSNR of the degraded image.
6 The distinct influences of early and late times on generation
The above experiments revealed striking differences of performance between models trained with different losses; naturally raising the question: why are some of the considered models much worse at generating than the FM baseline and at which stage of the generative process do they fail? A first step towards addressing this question is to develop a more fine-grained understanding of the temporal structure underlying the generation process.
6.1 Perturbating artificially the generation process
We first implement controlled perturbations of the denoiser, applied at different selected time intervals during the generation process.
More precisely, from the standard FM denoiser $D_{t}$ (trained with $\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$) we create controllable perturbed denoisers, equal to $D_{t}$ everywhere except on a given time interval $t_{\min},t_{\max}]$, where they are equal to $\tilde{D}_{t}\triangleq D_{t}+\sigma(t)\delta$. The controllable factors are:
  1. 1.
    the perturbation interval $t_{\min},t_{\max}]$. We consider intervals of length 0.3;
  2. 2.
    the level of perturbation $\sigma(t)$. We set it such that $\tilde{D}_{t}$ has a PSNR equal to $90\%$ of that of $D_{t}$;
  3. 3.
    the (deterministic) perturbation direction $\delta$. We consider (a) checkerboard perturbations corresponding to alternated patches of +1+1 and 1-1, with different patch sizes; (b) positive (resp. negative) shift respectively referring to a constant perturbation of +1+1 (resp. 1-1); (c) “Residual” referring to the perturbation $\tilde{D}_{t}:=D_{t}+\sigma(t)(\mathrm{Id}-D_{t})$ which can be seen as a relaxed denoiser. Perturbations are displayed in Appendix F, Figure 10.
Figure 4: Influence of different perturbations at different generation phases on the FID. Two classes of perturbations emerge: noise-type perturbations (checkboard 4x4, residual) characterized by high FID, low pairwise distance and strongest impact in the last times and drift-type perturbations (pos./neg. shift, checkerboard 16x16) characterized by low FID, high pairwise distance and strongest impact in the early times.
Results are presented Figure 4. The left plot displays the average $\ell_{2}$ distance between $500$ generated samples and their corresponding baseline samples (i.e., generated with $D_{t}$ starting from the same noise instance $x_{0}$). This allows to measure how the injected noise deviates the ODE trajectory and affects the generated sample. The right plot displays the test FID-10k for each perturbation applied on each interval.
We observe that the $1\times 1$ checkerboard perturbation does not induce any significative change in the generated samples, meaning it is effectively corrected after the perturbed time interval (although, by design of the experiment, like all perturbations it degrades the PSNR by 10 %). As the size of the patches grows from $1\times 1$ to $16\times 16$, the generated images deviate further from the initial samples. Something surprising happens: although larger patches always induce higher drift in the distribution (as assessed by the pairwise distance), we observe that applying the $4\times 4$ kernel size at later times produces a large increase in FID, while for the other checkboard perturbations, they remain quite low. Regarding the constant perturbations, they strongly alter the generated images, producing washed-out outputs (see Appendix F, Figures 11 and 12) but yield only a small FID change. On the contrary, the residual perturbation leads to a strong increase of the FID, although the generated images are noisy versions, close to the initial ones in $\ell_{2}$ distance and visually.
These findings lead us to the following takeaways:
  1. 1.
    First, despite all perturbations having, by design, the same impact on PSNR degradation, they do not affect FID in the same manner. A first remark is that it is possible to build denoisers with degraded denoising performance that still remain good generators. Second, the experiment shows a stronger sensitivity to noise-type perturbations (i.e. low patch-size checkerboard pattern or residual, both being Gaussian-like) than to drift-type ones (i.e. positive shift, negative shift, large patch-size checkerboard).
  2. 2.
    Second, drift-type perturbations, which induce global change to the full image, are most impactful when applied early, while noise-type perturbations, which act locally, have a stronger effect when applied in late time intervals. This aligns with previous studies on the effective receptive field of the velocity U-Net during generation: Kamb and Ganguli (2025, Fig. 4.a) and Niedoba et al. (2025, Fig. 3) show that it evolves from a large kernel that encompasses the full image (enabling the model to compute the average of the dataset at $t=0$, matching the closed-form target) to a local field (a few pixels) for the last time steps, corresponding to removal of very small noise.
(a) Splitting trajectories in 1D flow matching between a uniform $p_{0}$ and a discrete target $\hat{p}_{1}$, composed of 3 data points. Red arrows represent the optimal velocity $\hat{v}^{\star}(x_{t},t)$.
(b) Mean and standard deviation of the spectral norm $\|\nabla_{x}v(x_{t},t)\|_{2}$ computed on 1000 ODE trajectories on Cifar10 with Gaussian $p_{0}$.
Figure 5: The closed-form velocity shows a transition when trajectories split, producing an early peak in the Lipschitz constant. The trained models do not exhibit such distinction, and learn a smoother field.
6.2 Regularity discrepancies between target and learned models
Previous experiments showed the influence of the early phase of the generative process, sensitive to drift-type perturbations, and late phase, sensitive to noise-type perturbations. This distinction of small and large times has already been observed in different contexts: Biroli et al. (2024) identify different temporal regimes in the generative process, evolving from trajectories that are indistinguishable to trajectories that all converge toward the training dataset, thereby revealing a phase transition between generalisation and memorisation. Bertrand et al. (2025) show that learned velocities differ the most from the loss minimizer $\hat{v}^{\star}$ (4) at small and large $t$, and that small $t$ seem to be more important for creating new images.
To deepen our understanding of the various phases, we consider a simpler setup with $p_{0}$ the uniform distribution on $1,1]^{d}$, and the discrete target distribution $\hat{p}_{1}$. In this setting, the optimal velocity field $\hat{v}^{\star}$ admits a simple closed-form expression (Bertrand et al., 2025, Prop. 1). It is defined on cones  [^4] For a Gaussian $p_{0}$, the optimal velocity $\hat{v}^{\star}_{t}$ is defined on the whole space: the regions not covered by cones in the uniform case correspond to regions of very low probability in the Gaussian setting. $C^{(i)}=\{x\in\mathbb{R}^{d}:\exists t\in[0,1],x_{0}\in[-1,1]^{d},\,x=(1-t)x_{0}+tx^{(i)}\}$. The optimal velocity $\hat{v}^{\star}(x_{t},t)$ simply points towards the mean of training points $x^{(i)}$ for all the indices $i$ such that $x_{t}\in C^{(i)}$ (Bertrand et al., 2025, Prop. 1). It follows that for $t$ close to 0, $\hat{v}^{\star}(x_{t},t)$ points towards the mean of the dataset, whereas in the later steps, $x_{t}$ belongs to a single cone and $\hat{v}^{\star}(x_{t},t)$ points towards the associated data point (Figure 5(a)). In between, there is a transition phase when the trajectories “split” into different cones.
This simple setting examplifies the two phases of the target velocity: during small times, it rapidly changes; then, after some threshold $\tau$, points $x$ only belong to a single cone $\mathcal{C}^{(i)}$ and the velocity, equal to $\frac{x^{(i)}-x}{1-t}$, varies very smoothly. We conjecture that approximating the closed-form is the hardest at such critical $\tau$ because of a high local Lipschitz constant of the target velocity field with respect to $x$, making it difficult to accurately capture the trajectory splitting dynamics.
To test this hypothesis numerically, on Figure 5(b), we estimate the local Lipschitz constant of the velocity in $x$ at time $t\in[0,1]$ by computing the spectral norm of the Jacobian $\nabla_{x}v(x_{t},t)$ along sampled ODE trajectories, using the power method, both for the closed-form $\hat{v}^{\star}$ and for velocities induced by our set of denoisers. This quantity reflects how strongly trajectories diverge locally. Once again, we identify distinct temporal behaviours. First, in an early regime ($t\in[0.1,0.2]$), the closed-form velocity exhibits a sharp Lipschitz peak, when trajectories split into cones, which confirms our intuition. Our trained models, by contrast, fail to reproduce this peak. The pronounced gap – at a phase already identified as critical for generalisation (Bertrand et al., 2025) – shows that networks learn a smoother velocity, unable to match the closed-form, and that this actually helps trajectories drift away from training samples thereby favoring generalisation. Second, in an intermediate regime ($t\in[0.3,0.8]$), the models maintain a relatively high Jacobian spectral norm, consistently above the closed-form, with the best-performing models showing the largest values in this range. This shows that maintaining a high Lipschitz constant during intermediate times is not problematic; instead, it may reflect the capacity of neural networks to represent complex transformations (Salmona et al., 2022).
While only early times are critical when considering the closed-form (time when ODE trajectories split, inducing a peak in the Lipschitz constant), our experiment suggests that, when it comes to learned models, additional factors governing good generation may still occur later. We explore this intermediate phase further in the next section.
6.3 Probing the temporal phases with ad-hoc denoisers
In Section 6.1, we investigated how to artificially perturb a pretrained denoiser, identifying two types of behaviors: early-time drifts and late-time FID degradation. We now ask whether these effects can be reproduced directly through training, by modifying either the loss function or the class of functions over which the loss is minimized, while remaining within the framework of our denoising toolkit.
In order to explore the importance of matching the Lipschitz peak at $t=0.2$ of the closed form (see Figure 5(b)), we build a denoiser whose Jacobian spectral norm is softly penalized during training on early times (e.g. $[0.1,0.3]$). We fix as setup loss $\mathcal{L}_{\text{FM}}$, parametrization $\mathcal{C}_{\mathrm{I+NN}}$ and additional regularization denoted $\mathcal{R}_{\text{[0.1,0.3]}}$. Interestingly, this model matches the best performing model (standard FM) in FID with a Jacobian spectral norm that is twice the difference in interval $[0.1,0.3]$ (see Figures 14 and 15). This suggests that the early critical time identified for the closed-form is actually not the only decisive time when considering learned models. Conducting the mirror experiment, with regularization applied at late times, e.g. $\mathcal{R}_{\text{[0.5-0.8]}}$ applied on the interval $[0.5,0.8]$, leads to degraded generation performance. More precisely, while regularizing at early times produces models with similar FID than the FM baseline but that can generate visually different samples (see Figure 18), applying regularization at late times instead produces samples that are closer to the baseline models but with degraded FID, confirming the behaviors observed under artificial perturbations. We push further these experiments on regularized models in Section H.1.
In the same spirit of dissecting which temporal regions of the trajectory matter most, we now train models with ad-hoc loss weightings that deliberately bias learning toward specific times. We build a new denoiser with weighting $w_{t}^{\mathrm{mid}}=\frac{1}{(0.5-t)^{2}}$, putting the emphasis on accurate denoising at $t=0.5$ whereas traditional weights focus on small and large $t$ (Kim et al., 2025, Fig. 2). We test and analyze other mid weightings in Section H.2. Figure 6, displaying the distance between generated samples starting from the same noise, shows that $w_{t}^{\text{mid}}$ induces substantial deviations from both FM/den models, while generating points that are roughly at the same distance of the dataset, and having an FID similar to $\mathcal{L}_{\mathrm{den}}$ (full evaluation is in Appendix C). This model strongly differs from the others (see also generated samples in Figures 13 and 18), an interesting fact as several works suggest, on the contrary, that all models, irrespective of architecture and optimization (Niedoba et al., 2025), and even subset of training data (Kadkhodaie et al., 2024) end up generating the same data. [Uncaptioned image] Figure 6: Left: average pairwise distance inter-models computed on $500$ samples, sharing same $x_{0}$ across models (CelebA-64). Right: average distance of samples to train set. $\mathcal{L}_{\mathrm{mid}}$ produces samples that differ from those of other models.
It follows from this study that:
  1. 1.
    Similar FIDs can hide different generation behaviours. Our general denoiser framework makes it possible to explicitly build such models. By acting on the early/mid time of the generation process, we are able to change the samples generated.
  2. 2.
    We show a gap between the closed-form temporal properties (early ODE trajectories splitting vs. denoising with final image already determined) and the behavior of learned generative models, where the intermediate regime matters more.
7 Conclusion
While flow matching can in principle be equivalently recast as a denoising task, we showed that connecting them also reveals how alternative choices lead to substantial variations in model behaviour. Overall, our experiments reveal that the relationship between denoising accuracy and generative performance is more subtle and complex than it may appear: it is possible to construct denoisers with degraded denoising performance without affecting the FID. Our analysis shows that engineered perturbations affect models differently depending on when they occur: drift-type perturbations are most impactful early, while noise-type ones mostly impact later stages of the process. Comparing regularity dynamics of target and learned velocities further confirms this temporal asymmetry and shows the importance of intermediate times for generation with learned models. Incidentally, we observe that models with similar FID scores can nonetheless display distinct generative behaviours. Looking ahead, the importance of parametrization choices ($\mathcal{C}_{\mathrm{NN}}$ vs. $\mathcal{C}_{\mathrm{I+NN}}$), emphasized in our results, deserves further exploration – for instance through gradient-step denoisers. To facilitate such investigations, we will release our code and complete toolbox of trained models, providing the community with a resource to probe generative models beyond standard benchmarks.
Appendix A Details on the losses
Flow matching denoising loss
Substituting the velocity $v(x,t)=\tfrac{D(x,t)-x}{1-t}$ into the standard Flow Matching loss (Eq. (2)) together with $x_{1}-x_{0}=\frac{x_{1}-x_{t}}{1-t}$ yields
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{0}\sim p_{0},x_{1}\sim p_{1}\end{subarray}}\left[\|{v_{t}(x_{t})}-(x_{1}-x_{0})\|^{2}\right]$ $\quad=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{0}\sim p_{0},x_{1}\sim p_{1}\end{subarray}}\left[\left\|\frac{{D_{t}(x_{t})}-x_{t}}{1-t}-(x_{1}-x_{0})\right\|^{2}\right]$
$\quad=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{0}\sim p_{0},x_{1}\sim p_{1}\end{subarray}}\left[\left\|\frac{{D_{t}(x_{t})}-x_{t}-(x_{1}-x_{t})}{1-t}\right\|^{2}\right]$
$\quad={\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{0}\sim p_{0},x_{1}\sim p_{1}\end{subarray}}\left[\tfrac{1}{(1-t)^{2}}\,\|{D_{t}(x_{t})}-x_{1}\|^{2}\right]}$
$=\mathcal{L}_{\text{FM}}(D)$
Classical denoising loss.
We now compare the setting of generative denoisers that take $x_{t}=(1-t)x_{0}+tx_{1}$ as input with that of classical denoisers that take
$x_{\sigma}=x_{1}+\sigma x_{0}$
as input. A classical denoiser $\tilde{D}$ is usually trained by minimizing
$\mathcal{L}(\tilde{D})=\mathbb{E}_{\begin{subarray}{c}\sigma\sim\mathcal{U}([0,\sigma_{\max}])\\ x_{0}\sim p_{0},x_{1}\sim p_{1}\end{subarray}}\left[\|\tilde{D}(x_{\sigma},\sigma)-x_{1}\|^{2}\right],$ (13)
where $\sigma_{\max}$ is typically around $0.5$. From Remark 1, classical and generative denoisers are equivalent up to the reparameterization
$t(\sigma)=\frac{1}{1+\sigma}.$
A change of variables in (13) gives:
[…]
$=\mathbb{E}_{\begin{subarray}{c}x_{0}\sim p_{0}\\ x_{1}\sim p_{1}\end{subarray}}\left[\int_{0}^{1}\|\tilde{D}(x_{t}/t,(1-t)/t)-x_{1}\|^{2}\mathbf{1}_{[1/(1+\sigma_{\text{max}}),1]}(t)\frac{1}{t^{2}}\mathrm{d}t\right]$
$=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([(1+\sigma_{\text{max}})^{-1},1])\\ x_{0}\sim p_{0},x_{1}\sim p_{1}\end{subarray}}\left[\frac{1}{t^{2}}\|D(x_{t},t)-x_{1}\|^{2}\right],$
$=\mathcal{L}_{\text{classic}}(D),$
where in the before last line we used $D(x_{t},t)=\tilde{D}\!\left(\frac{x_{t}}{t},\frac{1-t}{t}\right)$.
Appendix B Details on training generation and metrics
All networks are trained with the same random initialization to ensure comparability. For CIFAR-10 we train for 400 epochs with batch size $128$, and for CelebA-64 we train for 300 epochs with batch size $128$. We apply exponential moving average to stabilize training. For CelebA-64, we use a U-Net architecture [Ronneberger et al., 2015] as in Ho et al. [2020]. For CIFAR-10, we adopt the architecture from the torchcfm library [Tong et al., 2024].
The 10 denoisers of Figure 2 are trained independently with 400 epochs each. All samples are generated by solving the ODE using the dopri5 scheme from $t=0$ to $t=1$ with $100$ timesteps; generating samples using a denoiser $D$ means that using the velocity $v(x,t)=\tfrac{D(x,t)-x}{1-t}$.
To measure generation quality, we use the standard Fréchet Inception Distance [Heusel et al., 2017] with the Inception-V3 features [Szegedy et al., 2016]; as recommended by Stein et al. [2023] we also compute it with the DINOv2 embedding.
Appendix C Additional results on CIFAR10 and CelebA-64
We report additional results on CIFAR10 and CelebA-64. In particular, we broaden the range of weights considered by including those defined in Table 2, with weights putting more or less emphasis than $w_{t}^{\text{FM}}$ on large times ($w_{t}^{\text{pow 1}}$, $w_{t}^{\text{pow 3}}$) and one putting more emphasis on $t=0.2$, $w_{t}^{\text{mid-0.2}}$.
Tables 3 and 4 present the performance of all models in terms of PSNR and the two geenrative metrics, FID and DINO. Figure 7 provides a visual overview for CelebA-64, as was done for CIFAR-10 in the main paper.
Table 2: Summary of the denoising losses (left) and parametrization classes (right).
Losses $\mathcal{L}$ Parametrization classes $\mathcal{C}$
$\mathcal{L}_{\text{FM}}:\,w_{t}^{\text{FM}}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{NN}}$: $D(x,t)=N^{\theta}(x,t)$
$\mathcal{L}_{\text{classic}}:\,w_{t}^{\text{classic}}=\frac{1}{t^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$: $D(x,t)=x+(1-t)N^{\theta}(x,t)$
$\mathcal{L}_{\text{den}}:\,w_{t}^{\text{den}}=1$
$\mathcal{L}_{\text{pow 1}}:\,w_{t}^{\text{pow 1}}=\frac{1}{(1-t)}$
$\mathcal{L}_{\text{pow 3}}:\,w_{t}^{\text{pow 3}}=\frac{1}{(1-t)^{3}}$
$\mathcal{L}_{\text{mid}}:\,w_{t}^{\text{mid}}=\frac{1}{(0.5-t)^{2}}$
$\mathcal{L}_{\text{mid-0.2}}:\,w_{t}^{\text{mid-0.2}}=\frac{1}{(0.2-t)^{2}}$
Table 3: PSNR and FID for the different losses, to be compared with the standard FM (bottom line, corresponding to loss $\mathcal{L}_{\text{FM}}$ and parametrization $\mathcal{C}_{\mathrm{I+NN}}$) . PSNR computed on 1000 images; FID on 10k test images; CIFAR-10, 400 epochs.
Loss Class PSNR (↑) FID / DINO (↓)
$t=0.1$ $t=0.3$ $t=0.6$ $t=0.9$ $t=0.95$
$\mathcal{L}_{\text{FM}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 14.41 18.16 23.42 32.89 37.33 12.97 / 506.54
$\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\text{NN}}$ 14.42 18.17 23.35 32.54 36.70 26.72 / 721.33
$\mathcal{L}_{\text{classic}}$ $w_{t}=\frac{1}{t^{2}}\mathbf{1}_{t>t_{\min}}$ $\mathcal{C}_{\text{NN}}$ 14.41 17.99 22.74 30.27 32.29 101.12 / 1444.69
$\mathcal{L}_{\text{pow 1}}$ $w_{t}=\frac{1}{(1-t)}$ $\mathcal{C}_{\text{NN}},$ 14.42 18.20 23.44 32.80 37.17 12.13 / 459.96
$\mathcal{L}_{\text{pow 3}}$ $w_{t}=\frac{1}{(1-t)^{3}}$ $\mathcal{C}_{\text{NN}}$ 14.33 17.96 23.14 32.72 37.23 45.15 / 1189.55
$\mathcal{L}_{\text{mid}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 13.91 18.05 23.51 32.85 37.16 22.34 / 527.53
$\mathcal{L}_{\text{mid-0.2}}$ $w_{t}=\frac{1}{(0.2-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 14.34 18.20 23.41 32.66 36.90 22.67 / 611.83
$\mathcal{L}_{\text{den}}$ $w_{t}=\frac{1}{(1-t)}$ $\mathcal{C}_{\mathrm{I+NN}},$ 14.42 18.18 23.35 32.58 36.79 19.29 / 717.76
$\mathcal{L}_{\text{classic}}$ $w_{t}=\frac{1}{(1-t)}$ $\mathcal{C}_{\mathrm{I+NN}},$ 14.42 17.99 22.78 30.53 34.09 148.95 / 1681.69
$\mathcal{L}_{\text{pow 1}}$ $w_{t}=\frac{1}{(1-t)}$ $\mathcal{C}_{\mathrm{I+NN}},$ 14.42 18.20 23.44 32.80 37.17 12.64 / 481.42
$\mathcal{L}_{\text{pow 3}}$ $w_{t}=\frac{1}{(1-t)^{3}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.39 18.13 23.36 32.90 37.42 15.32 / 564.80
$\mathcal{L}_{\text{mid}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 13.93 18.04 23.51 32.86 37.22 17.34 / 564.80
$\mathcal{L}_{\text{FM}}+\mathcal{R}_{\text{early}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.18 18.19 23.52 32.97 37.43 10.34 / 323.22
$\mathcal{L}_{\text{FM}}+\mathcal{R}_{\text{late}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.41 17.92 23.43 32.97 37.43 22.81 / 565.44
10-denoisers, $\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\mathrm{NN}}$ 14.42 18.16 23.39 33.02 37.49 9.80 / 265.04
10-denoisers $\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.42 18.18 23.40 33.03 37.50 9.44 / 248.00
$\mathcal{L}_{\text{FM}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.41 18.22 23.52 32.98 37.44 9.64 / 303.03
[…]
Loss Class PSNR (↑) FID / DINO (↓)
$t=0.1$ $t=0.3$ $t=0.6$ $t=0.9$ $t=0.95$
$\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\text{NN}}$ 16.01 20.98 26.25 34.50 37.91 21.13 / 511.97
$\mathcal{L}_{\text{classic}}$ $w_{t}=\frac{1}{t^{2}}\mathbf{1}_{t>t_{\min}}$ $\mathcal{C}_{\text{NN}}$ 15.88 20.19 24.41 29.35 29.81 87.80 / 1706.42
$\mathcal{L}_{\text{pow 1}}$ $w_{t}=\frac{1}{(1-t)}$ $\mathcal{C}_{\text{NN}}$ 16.04 21.07 26.44 34.99 38.84 8.36 / 271.09
$\mathcal{L}_{\text{pow 3}}$ $w_{t}=\frac{1}{(1-t)^{3}}$ $\mathcal{C}_{\text{NN}}$ 15.51 20.39 25.79 34.75 38.86 34.39 / 751.94
$\mathcal{L}_{\text{mid}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 15.29 20.82 26.54 35.17 38.95 22.25 / 441.85
$\mathcal{L}_{\text{mid-0.2}}$ $w_{t}=\frac{1}{(0.2-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 15.98 21.12 26.49 34.93 38.59 11.10 / 356.50
$\mathcal{L}_{\text{FM}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\text{NN}}$ 15.93 20.79 26.12 34.83 38.81 15.48 / 444.89
$\mathcal{L}_{\text{den}}$ $w_{t}=1$ $\mathcal{C}_{\mathrm{I+NN}}$ 15.98 20.87 26.10 34.40 38.06 19.02 / 508.06
$\mathcal{L}_{\text{classic}}$ $w_{t}=\frac{1}{t^{2}}\mathbf{1}_{t>t_{\min}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 15.89 20.26 24.70 31.34 34.67 135.64 / 1368.35
$\mathcal{L}_{\text{pow 1}}$ $w_{t}=\frac{1}{(1-t)}$ $\mathcal{C}_{\mathrm{I+NN}}$ 16.05 21.07 26.46 35.09 39.03 6.93 / 248.00
$\mathcal{L}_{\text{pow 3}}$ $w_{t}=\frac{1}{(1-t)^{3}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 15.98 20.98 26.40 35.28 39.39 7.10 / 209.17
$\mathcal{L}_{\text{mid}}$ $w_{t}=\frac{1}{(0.5-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 14.80 20.79 26.58 35.22 39.08 21.17 / 419.67
$\mathcal{L}_{\text{FM}}$ $w_{t}=\frac{1}{(1-t)^{2}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 16.03 21.08 26.52 35.33 39.34 5.33 / 167.68
(a) Difference in PSNR (lower is better) between standard FM ($\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$) and various models, computed on 1000 test images. Positive values indicate worse denoising performance compared to standard FM.
(b) FID on 10k tests images.
Figure 7: PSNR and FID for the different losses and parametrizations, CelebA-64, 300 epochs. Models that reach the best PSNR (low values of PSNR difference with respect to standard FM) also reach the lowest FID.
Appendix D Sensitivity of generation to last timesteps
Figure 8: FID comparison when replacing late-time denoising with a pre-trained GS-DRUNet. While the external denoiser slightly improves performance, a gap with the FM baseline remains, suggesting late-time denoising is not the main cause of the discrepancy.
Given the sensitivity of FID to noise, one might wonder whether the poor performance of some denoisers in our toolkit is due to their limited denoising ability at late times (i.e., low noise levels). The following experiment LABEL:{fig:fid_drunet} shows that this is not the case. We generate samples as usual with a Dopri5 scheme, but stop the integration at $t=0.95$ and complete the process with a generic GS-DRUNet pre-trained denoiser from the deepinv library [Tachella et al., 2025], trained on a different dataset. We then compute the FID on 10k generated images and compare it with the original FID. Although plugging in this external denoiser slightly improves the FID, a gap with the FM baseline remains, indicating that late-time denoising is not the only cause of the discrepancy.
Appendix E Inpainting visual results
We display in Figure 9 the reconstructions obtained with our denoisers on a single inpainting task on CelebA-64 with a mask of size $17\times 17$. We set the parameter $\alpha$ in PnP-Flow to $\alpha=0.3$ and the number of iterations to $100$, following the recommendations of the original paper [Martin et al., 2025] and using their public implementation. One can see that only a few methods accurately recover the headband: namely the flow-matching baseline $\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$ and the variants $\mathcal{L}_{\text{mid}},\mathcal{C}_{\text{NN}}$ and $\mathcal{L}_{\text{mid}},\mathcal{C}_{\mathrm{I+NN}}$.
(a) Noisy
(b) $\mathcal{L}_{\text{den}},\mathcal{C}_{\text{NN}}$
(c) $\mathcal{L}_{\text{classic}},\mathcal{C}_{\text{NN}}$
(d) $\mathcal{L}_{\text{pow\_1}},\mathcal{C}_{\text{NN}}$
(e) $\mathcal{L}_{\text{pow\_3}},\mathcal{C}_{\text{NN}}$
(f) $\mathcal{L}_{\text{mid}},\mathcal{C}_{\text{NN}}$
(g) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\text{NN}}$
(h) $\mathcal{L}_{\text{den}},\mathcal{C}_{\mathrm{I+NN}}$
(i) $\mathcal{L}_{\text{classic}},\mathcal{C}_{\mathrm{I+NN}}$
(j) $\mathcal{L}_{\text{pow\_1}},\mathcal{C}_{\mathrm{I+NN}}$
(k) $\mathcal{L}_{\text{pow\_3}},\mathcal{C}_{\mathrm{I+NN}}$
(l) $\mathcal{L}_{\text{mid}},\mathcal{C}_{\mathrm{I+NN}}$
(m) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$
Figure 9: Inpainting results: top row $\mathcal{C}_{\text{NN}}$, bottom row $\mathcal{C}_{\mathrm{I+NN}}$, columns correspond to losses.
Appendix F Perturbation samples
Figure 10: Perturbations applied to denoiser $D_{t}$ (here for $t=0.3)$. Experiments done on CIFAR-10.
Figure 11: Unperturbed images.
Interval Ckb. 1x1 Ckb. 4x4 Ckb. 8x8 Ckb. 16x16 Pos. shift Neg. shift Residual
$[0.,0.3]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.3,0.6]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.6,0.9]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.,0.3]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.3,0.6]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.6,0.9]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Effect of perturbations on generated samples (CIFAR-10).
Appendix G Samples obtained on CelebA-64
(a) $\mathcal{L}_{\text{den}},\mathcal{C}_{\text{NN}}$
(b) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\text{NN}}$
(c) $\mathcal{L}_{\text{den}},\mathcal{C}_{\mathrm{I+NN}}$
(d) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$
(e) $\mathcal{L}_{\text{mid-0.2}},\mathcal{C}_{\mathrm{NN}}$
(f) $\mathcal{L}_{\text{mid}},\mathcal{C}_{\text{NN}}$
(g) $\mathcal{L}_{\text{den}},\mathcal{C}_{\text{NN}}$
(h) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\text{NN}}$
(i) $\mathcal{L}_{\text{den}},\mathcal{C}_{\mathrm{I+NN}}$
(j) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$
(k) $\mathcal{L}_{\text{mid-0.2}},\mathcal{C}_{\mathrm{NN}}$
(l) $\mathcal{L}_{\text{mid}},\mathcal{C}_{\text{NN}}$
Figure 13: Sampling results on CelebA-64 for different models and two initial points $x_{0}$ (shared across columns).
Appendix H Additional experiments on generation phases
H.1 Time-interval Jacobian penalized models
Experimental setup
We train standard flow matching models ($\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$) with additional Jacobian spectral norm regularization applied over a prescribed time interval $t_{\mathrm{min}},t_{\mathrm{max}}]$:
$\mathcal{R}_{[t_{\mathrm{min}},t_{\mathrm{max}}]}(\theta):=\lambda\mathbf{1}_{t\in[t_{\mathrm{min}},t_{\mathrm{max}}]}\max\left(\|\nabla_{x}v^{\theta}_{t}(x_{t})\|_{2},M\right),$ (14)
where $\lambda$ is a regularization parameter and $M$ is the targeted upper bound on the Jacobian spectral norm. The models are trained without regularization for $390$ epochs and finetuned with regularization for the last $10$ epochs. The Jacobian spectral norm is estimated using the power method ($10$ iterations). The threshold $M$ is set to $M=4$ for early intervals (i.e. $[0,0.2],[0.1,0.3]$) and $M=2$ for the others. The regularization parameter is set to $\lambda=0.1$ for all intervals except for $[0.3,0.6]$ and $[0.5,0.8]$ where a more aggressive regularization $\lambda=0.2$ is used to ensure an effective reduction of the Jacobian spectral norm on these intervals. Indeed, our goal here is to induce a controlled decrease of the Jacobian norm over selected time intervals and study how this affects denoising and generation.
Comparison with the closed-form optimal velocity field.
The early-time regularizations (i.e. $[0,0.2]$ and $[0.1,0.3]$) produce the largest deviation from the closed-form behaviour in terms of the mean Jacobian spectral norm measured over the ODE trajectories (see Figure 14(a)): models with these early-time regularizations fail to reproduce the sharp peak of the closed-form Jacobian around $t\approx 0.2$ and instead maintain higher values in mid times. In contrast, models regularized at mid or late times begin by increasing the Jacobian spectral norm (similarly to standard FM) and then exhibit a decay around mid times, getting closer to the closed-form behaviour.
Comparison with standard Flow Matching.
In terms of denoising performance (Figure 15(a)), a decrease of the Jacobian norm over a given interval (compared to the standard FM model) corresponds to a decrease in PSNR at time steps within that interval. In other words, constraining the local Lipschitz constant deteriorates the denoising accuracy at these times.
In terms of generation performance (Figure 15(b)), two classes of models clearly appear. Penalizing the Jacobian norm in early intervals ($[0,0.2]$, $[0.1,0.3]$) achieves FID scores comparable to standard FM. In contrast, penalizing it in the other intervals results in higher FIDs (especially for mid time intervals $[0.3,0.6]$, $[0.3,0.8]$). To better quantify the impact of these regularizations on generated samples, we compute the average pairwise distance between samples generated from the same $x_{0}$, thus directly comparing ODE trajectories (Figure 17(a)). Among all models, larger distances relative to the standard FM baseline are observed for penalizations applied at early and mid times. Qualitatively (see Figure 18), early-time penalization ($[0.1,0.3]$) can lead to visually distinct samples, sometimes even changing the image class, whereas late-time regularization mostly yields small perturbations or noisy versions compared to the standard FM samples.
This aligns with observations made from the controlled perturbation experiments in Section 6.1, where macro-level perturbations applied at early times tend to shift the global image distribution but maintain a low FID, whereas micro-level perturbations applied at late times produce noisier samples and lead to a stronger degradation of the FID.
Discussion
Overall, this experiment indicates that shifting the peak of the Jacobian spectral norm as done with the early-time penalizations can change the ODE trajectories without altering the generation quality. In contrast, maintaining a relatively high Jacobian norm appears important in order to keep good denoising quality in the mid times, which in turn is crucial to achieve good generation quality. Notably, matching the behaviour of the closed-form velocity field in terms of the Jacobian spectral norm seems undesirable in the early/mid-time regimes: not reproducing its peak at early times still yields models with competitive FID scores, whereas getting closer to the closed-form behaviour at mid times results in degraded FIDs. This can be understood from the perspective of the closed-form denoiser, which at mid times reproduces samples from the training dataset: distinct input points are mapped to the same output, i.e. image of the training set, which is not the expected behaviour of a good denoiser or a good generator.
H.2 Models with intermediate weighting
Experimental setup.
To dissect which temporal regions of the diffusion trajectory contribute most to the learned denoising dynamics, we extend our denoiser toolkit with a family of ad-hoc objectives that deliberately bias learning toward specific times. In particular, we introduce a weighting function
$w_{t}^{\mathrm{mid-}t^{*}}=\frac{1}{(t^{*}-t)^{2}},$ (15)
which emphasizes accurate denoising around a chosen time $t^{*}\in[0,1]$. This formulation generalizes the intermediate weighting $w_{t}^{\mathrm{mid}}$ presented in Section 6.2, allowing us to probe the model’s sensitivity to localized temporal emphasis along the diffusion process. As with the baseline models, we evaluate for each $w_{t}^{\mathrm{mid-}t^{*}}$ both the per-time PSNR curves (Figure 16(a)) and the final FID scores (Figure 16), complemented by pairwise model distance maps (Figure 17(b)) and the evolution of spatial regularity measured through the Jacobian spectral norm (Figure 14(b)).
Observations and discussion.
Surprisingly, despite the formal resemblance between the standard flow-matching weighting $w_{t}^{\mathrm{FM}}=\frac{1}{1-t}^{2}$ and the limiting behavior of $w_{t}^{\mathrm{mid-}t^{*}}$ as $t^{*}\to 1$, the results diverge markedly. Instead of recovering the performance of the FM baseline, we observe a sharp degradation: the FID explodes for $t^{*}=0.95$, and the PSNR deteriorates across all times, including $t=0.95$ itself. This finding indicates that it is not straightforward to isolate the influence of a single time point: poor denoising performance at other times propagates through the training dynamics, affecting even the regions of emphasis. Interestingly, the model most similar to the FM-shifted baseline is obtained for $t^{*}=0.2$. These effects are not attributable to the architectural choice of class $\mathcal{C}_{\mathrm{NN}}$ versus $\mathcal{C}_{\mathrm{I+NN}}$, since the pair $(\mathcal{L}_{\mathrm{FM}},\mathcal{C}_{\mathrm{NN}})$ also has a very good FID.
Two distinct degradation regimes emerge in the FID/PSNR trends, echoing the observations of Section 6.1. In the first phase, when the emphasis is placed on intermediate times $t^{*}\in[0.2,0.5]$, the FID remains roughly constant and close to that of the uniform weighting $w_{t}^{\mathrm{den}}=1$, yet the samples display visible drifts (Figure 19). These models appear to have learned qualitatively different denoising functions, offering insights into how weighting can modulate trajectory alignment. In the second phase, corresponding to larger times $t^{*}\in[0.6,0.9]$, image quality degrades sharply, both in terms of FID and PSNR.
Pairwise distance analysis corroborates these observations: the distance from samples generated using the baseline model increases monotonically with $t^{*}$, with a mild peak around $t^{*}=0.7$. We hypothesize that this pattern reflects two regimes: a “drift perturbation” regime for lower $t^{*}$, where structural shifts dominate, and a “noise amplification” regime for higher $t^{*}$, where instability in late-time dynamics prevails. Indeed, models with $t^{*}=0.5$ and $t^{*}=0.8$ exhibit comparable distances to the baseline while producing qualitatively distinct images (Figure 19).
Finally, these findings align with our analysis of the Lipschitz constant. Models that drastically reduce their Lipschitz constant relative to the FM baseline in the interval $t\in[0.3,0.6]$ coincide with those exhibiting catastrophic FID degradation. Conversely, “drift-perturbed” models tend to suppress the Lipschitz constant primarily for $t\in[0,0.3]$, while maintaining higher expressivity in the mid-range $t\in[0.3,0.6]$.
(a) Jacobian spectral norm for regularized models.
(b) Jacobian spectral norm for the different mid losses.
Figure 14: Mean and standard deviation of the spectral norm $\|\nabla_{x}v(x_{t},t)\|_{2}$, computed on 1000 ODE trajectories $(x_{t})$. CIFAR-10, 400 epochs. The Jacobian spectral norm is estimated using the power iteration with maximum number of iteration set to $10$.
(a) Difference in PSNR (lower is better) between standard FM ($\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$) and various models, computed on 1000 test images. Positive values indicate worse denoising performance compared to standard FM.
(b) FID on 10k tests images.
Figure 15: PSNR and FID for the different regularizations, CIFAR-10, 400 epochs. PSNR degradation at early times does not impair generation performance, while PSNR degradation at mid times systematically correlates with a higher FIDs.
(a) Difference in PSNR (lower is better) between standard FM ($\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$) and various models, computed on 1000 test images. Positive values indicate worse denoising performance compared to standard FM.
(b) FID on 10k tests images.
Figure 16: PSNR and FID for the different mid losses, CIFAR-10, 400 epochs.
(a) Average pairwise distance for the different Jacobian regularizations.
(b) Average pairwise distance for the different mid losses.
Figure 17: Average pairwise distance inter-models computed on $1000$ samples, sharing same $x_{0}$ across models. Experiments done on CIFAR-10, models trained with 400 epochs.
$\mathcal{L}_{\mathrm{FM}},\ \mathcal{C}_{\mathrm{I+NN}}$ $\mathcal{L}_{\mathrm{FM}}+\mathcal{R}_{0.1-0.3},\ \mathcal{C}_{\mathrm{I+NN}}$ $\mathcal{L}_{\mathrm{FM}}+\mathcal{R}_{0.5-0.8},\ \mathcal{C}_{\mathrm{I+NN}}$
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 18: CIFAR-10 samples generated by models trained with different Jacobian regularizations. Regularizing at early times ($\mathcal{R}_{0.1-0.3}$) changes visually more the samples than applying regularization at later times ($\mathcal{R}_{0.5-0.8}$).
$\mathcal{L}_{\mathrm{FM}},\ \mathcal{C}_{\mathrm{I+NN}}$ $\mathcal{L}_{\mathrm{mid0.5}},\ \mathcal{C}_{\mathrm{NN}}$ $\mathcal{L}_{\mathrm{mid0.8}},\ \mathcal{C}_{\mathrm{NN}}$
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 19: CIFAR-10 samples generated by models trained with different mid weightings. Each row corresponds to one fixed initial point $x_{0}$.